# Space-time first-order correlations of an open Bose Hubbard model with incoherent pump and loss

### v1

@@ Appendix D Steady-state solution for hard-core bosons
where the identity operator is the sum over the complete Fock basis for the $N$-particle sector. The number of states in the system of hard-core bosons is identical to the one of the fermionic system, which explains the identical normalisation $\mathcal{N}_{F}$ and structure of the solution. One can show that the Hamiltonian conserves the particle number, or equivalently the number of up-spins, and hence it commutes with the steady state density matrix. The dissipative part of the Lindbladian evolution vanishes with the same reasoning as for the bosons and fermions.
Appendix E Decay rate at large interactions
@@
For $J=0$, each site in the Keldysh action (2.3) decouples. Hence, in the strong interaction limit $U\rightarrow\infty$, one is left with solving the problem of $L$ independent dissipative Kerr resonators. This is a well-known exactly solvable problem [59]. One can exploit a weak symmetry to block-diagonalise the Lindbladian and then introduce an operator which renders the Lindbladian quadratic in the superoperators. Equivalently, one can introduce a time-dependent gauge transformation to obtain a quadratic action as shown in [50]. The Kerr non-linearity then transforms into a fluctuating frequency depending on the number density which in turn leads to dephasing. The result for the retarded Green’s function is given by [59]
@@
The last term is suppressed in the large $t$ limit. Expanding the second term for $U\gg J$, we find that
@@
such that the renormalised decay for $U\gg J$ reads
@@
Appendix F Details on the numerical calculations
F.1 Exact diagonalisation
Numerical results were obtained by exact diagonalisation of the Hamiltonian or Lindbladian in the Fock basis of particle occupation numbers per site and the time evolution implemented in the qutip library [73, 74]. In the case of the closed-system evolution, we additionally use the conservation of particle number, as explained in section F.5.
F.2 Fit of the decay
@@
Restricting the dimension of the local Hilbert space, we obtain results for all interaction strength by exact diagonalisation. Using this method, we extract the slope of ${\cal G}(t)$ by fitting a family of linear curves over different intervals. The multiple scale dependence of the oscillation of the two-point function makes this fitting procedure necessary. As the result depends on the chosen interval of the fit, we fit a family of 20 curves in the interval accessible in order to reach sufficient numerical precision. The average and variance are shown as an example in Fig. 8.
@@
Figure 8: Logarithm of the retarded Green’s function ${\cal G}(t)=-\Re\Big{[}\ln\text{i}G_{00}^{R}(0t)\Big{]}$ (dimensionless) obtained from exact diagonalisation (red) together with the single-pole approximation (black, solid) as a function of time (in units of $\tau_{l}$) for the site $j=0$. The corresponding fitted decay rate is indicated by the black solid line. The parameters are $U=5.623$, $J=1.0$, $z=0.1$, $L=4$, $N_{s}=4$.
F.3 Effect of the Hilbert space truncation
@@
We showed in the main text that we recover the analytically predicted renormalisation of the decay $\kappa(U)$ for small hopping $J$. In order to benchmark the numerical truncation over the local Hilbert space, we here investigate the dependence of $\kappa(U)$ for large $J=10$. We show in Fig. 9 the analytical (one-sited) prediction for small $J$ and the numerical results for $N_{s}\in\{3,4,5\}$. For large $U$, we expect a deviation from the one-sited result and we observe that the reached plateau is independent of the local Hilbert space truncation. In the weak interaction limit, we interpret the closeness of the results for $N_{s}=4$, $N_{s}=5$ with the exact prediction as an indication of numerical convergence.
@@
Figure 9: Decay $\kappa(U)$ for $J=10$, $z=0.1$ and $L=4$, and for different local truncation of the Hilbert space $N_{s}$ in the numerics. We expect to recover the non-interacting solution for small $U$ and a deviation from the one-sited solution in the limit of large interaction.
F.4 System size dependence
@@
We further investigate the dependence of the reached plateau value of the decay for large $U$ and $J$ on the system size. From Fig. 9, we conclude that a truncation of $N_{s}=3$ is sufficient to reach the converged plateau value. We hence fix $N_{s}=3$ and vary the system size $L$. The dependence of the renormalised decay on $L$ is shown in Fig. 10, in a limit of large hopping and interaction ($J=10$ and $U=100$). We observe that the effective decay increases monotonously with the system size. Further analytical insight could be obtained from the perturbative calculation of corrections in $J$ in the strong interaction limit around the exact result (58).
@@
Figure 10: Decay $\kappa(U)$ for different system sizes $L$. The parameters are $J=10$, $z=0.1$ and $N_{s}=3$.
F.5 Exact diagonalisation for the unitary system
@@
We calculate the Green’s function as the overlap of two time-evolved states. Since we are interested in the unitary time evolution, the dimensionality of the problem greatly reduces, and is further diminished by the symmetries present in the problem (particle number conservation, momentum conservation and parity conservation). We calculate the Green’s function as
@@
F.6 Simulations with matrix product states
In order to produce the plots in Fig. 6 we employ a representation of the many-body state as a matrix-product state [75] and code our algorithm using the Itensors library [76, 77]. We prepare the ground state considering a maximum number of bosons per site equal to $2$ on a lattice with periodic boundary condition and at density $N/L=1/4$. We subsequently compute the lesser and greater Green’s functions $G^{<}_{0j}(t)=-i\langle b_{j}^{\dagger}(t)b_{0}\rangle$ and $G^{>}_{0j}(t)=-i\langle b_{0}b_{j}^{\dagger}(t)\rangle$ and using these data we reconstruct the spectral functions plotted in the figure. For the time evolution, we fix $J=1$ and use a time step of $\delta t=0.01$.

### v2

@@ Appendix D Steady-state solution for hard-core bosons
where the identity operator is the sum over the complete Fock basis for the $N$-particle sector. The number of states in the system of hard-core bosons is identical to the one of the fermionic system, which explains the identical normalisation $\mathcal{N}_{F}$ and structure of the solution. One can show that the Hamiltonian conserves the particle number, or equivalently the number of up-spins, and hence it commutes with the steady-state density matrix. The dissipative part of the Lindbladian evolution vanishes with the same reasoning as for the bosons and fermions.
Appendix E Detailed calculations in Keldysh formalism
E.1 Fourier conventions
We first define the Fourier conventions used in this work. For a lattice of size $L=Na$, where $a$ is the lattice spacing and $N\in 2\mathbb{N}$, we define a set of integers =$\mathcal{I}=\{-N/2+1,\dots,N/2\}$, such that the discrete positions and discrete momenta are given by
$x_{j}=ja\qquad j\in\mathcal{I}\,,\qquad\qquad p_{k}=k\frac{2\pi}{L}\qquad k\in\mathcal{I}\,.$ (61)
We define the Fourier transforms by
$f(p_{k},\omega)$ $=\sum_{j\in\mathcal{I}}\int_{-\infty}^{\infty}\text{d}t\,f(x_{j},t)\,\text{e}^{-\text{i}p_{k}x_{j}}\,\text{e}^{\text{i}\omega t}\,\;,$ (62)
$f(x_{j},t)$ $=\sum_{k\in\mathcal{I}}\frac{a}{L}\int_{-\infty}^{\infty}\frac{\text{d}\omega}{2\pi}\,f(p_{k},\omega)\,\text{e}^{\text{i}p_{k}x_{j}}\,\text{e}^{-\text{i}\omega t}\,\,.$ (63)
In the thermodynamic limit $L\rightarrow\infty$, the momenta become continuous and the discrete sum tends to continuous integrals
$\sum_{k\in\mathcal{I}}\frac{1}{N}F(p_{k})\rightarrow a\int_{-\pi/a}^{\pi/a}\frac{\text{d}p}{2\pi}F(p)\,,$ (64)
for arbitrary functions $F$. Note that, in the main text, as well as in the numerics, we set $a=1$.
E.2 Dispersion line and spectral function
We calculate the inverse bare propagator as the second functional derivative of the action evaluated at vanishing field expectation values. Gathering the fields in a vector $\varphi=(\varphi_{c},\varphi_{c}^{*},\varphi_{q},\varphi_{q}^{*})^{T}$, this yields
$\Big{[}G_{0,m\ell}^{-1}\Big{]}_{\alpha\beta}(t,t^{\prime})\equiv\frac{\delta^{2}S[\varphi]}{\delta\varphi_{\alpha,m}(t)\delta\varphi_{\beta,\ell}(t^{\prime})}\,\Big{|}_{\varphi_{c,q}=0}\equiv\begin{pmatrix}{\mathbb{0}}_{2\times 2}&&{\mathbb{P}}^{A}_{m\ell}(t,t^{\prime})\\ {\mathbb{P}}^{R}_{m\ell}(t,t^{\prime})&&{\mathbb{P}}^{K}_{m\ell}(t,t^{\prime})\end{pmatrix}\;,$ (65)
where the subscripts $m,\ell$ label the discrete position indices and $\alpha,\beta$ the field components. The 0 index in $G_{0}$ indicates that this is the ‘bare’ propagator. Since it is evaluated at zero fields, it corresponds to the quadratic (non-interacting) theory. We obtain the inverse Keldysh component as ${\mathbb{P}}_{m\ell}^{K}(t,t^{\prime})=\text{i}\gamma\sigma_{x}\delta_{m\ell}\delta(t-t^{\prime})$, and the inverse of the retarded sector of the propagator as
${\mathbb{P}}^{R}_{m\ell}(t,t^{\prime})$ $=\begin{pmatrix}0&&-(\text{i}\partial_{t}+\text{i}\kappa_{0})\delta_{m,\ell}+J_{m,\ell}\\ (\text{i}\partial_{t}+\text{i}\kappa_{0})\delta_{m,\ell}+J_{m,\ell}&&0\end{pmatrix}\delta(t-t^{\prime})\;,$ (66)
where $J_{m\ell}$ is a short-hand notation for $J_{m\ell}\equiv(\delta_{m,\ell+1}+\delta_{m,\ell-1})$, and we defined the bare decay rate as $\kappa_{0}=({\gamma_{l}-\gamma_{p}})/{2}$ and $\gamma=\gamma_{l}+\gamma_{p}$. In order to calculate the excitation branches $\Omega(k)$ in this approximation, we search for the solutions implicitly given by [54, 79]
$\det{\mathbb{P}}^{R}(k,\Omega(k))=0\;,$ (67)
where ${\mathbb{P}}^{R}(k,\omega)$ is the Fourier transform in space and time of ${\mathbb{P}}^{R}_{m\ell}(t,t^{\prime})$ defined according to (62). Note that it only depends on one momentum and one frequency because of translational invariance in space and time. From (66) we find
$\Omega_{\pm}(k)=-\text{i}\kappa_{0}\mp 2J\cos(ka)\;,$ (68)
which is the solution given in the main text with a lattice spacing $a=1$.
E.3 Occupation number
The average occupation number $\langle n_{j}\rangle$ is encoded, within the Keldysh formalism, in the Keldysh Green’s function $G^{K}$, which is related to the second quantization operators by
$G_{jj}^{K}(t,t)=-\text{i}\Big{\langle}\{b_{j}(t),b^{\dagger}_{j}(t)\}\Big{\rangle}=-\text{i}(2\langle n_{j}\rangle+1)\,.$ (69)
By inverting the inverse bare propagator $G_{0}^{-1}$, one obtains the bare Keldysh component $G_{0}^{K}$ defined in (75) as
$G_{jj,0}^{K}(t,t)=a\int_{-\pi/a}^{\pi/a}\frac{\text{d}p}{2\pi}\int_{-\infty}^{\infty}\frac{\text{d}\omega}{2\pi}\frac{-\text{i}\gamma}{(\omega+2J\cos(pa))^{2}+\kappa_{0}^{2}}=-\text{i}\dfrac{\gamma}{2\kappa_{0}}=-\text{i}\dfrac{\gamma_{l}+\gamma_{p}}{\gamma_{l}-\gamma_{p}}\;,$ (70)
from which one deduces
$\langle n_{j}\rangle\equiv\langle n\rangle=\frac{1}{2}\Big{(}\frac{\gamma_{l}+\gamma_{p}}{\gamma_{l}-\gamma_{p}}-1\Big{)}=\frac{z}{1-z}\;,$ (71)
which does not depend on space as expected for a homogeneous system. This coincides with the result (12).
E.4 Perturbative corrections
We now turn to the calculation of the one-loop correction to the self-energy defined in (20). First, we introduce into the partition function $Z$ a source term $J=(J_{q}^{*},J_{q},J_{c}^{*},J_{c})^{T}$ linearly coupled to the field $\varphi$ as
$Z[J]=\text{e}^{\text{i}W[J]}\,=\int\mathcal{D}\varphi\;\exp\Bigg{\{}\text{i}S[\varphi]+\text{i}\int\text{d}t\sum_{j}\varphi_{j}^{T}(t)J_{j}(t)\Bigg{\}}\;.$ (72)
The Schwinger functional $W[J]$ is the generating functional of $n$-point connected correlation functions $G^{(n)}$, which are defined as
$G^{(n)}_{a_{1}\dots a_{n}}\equiv\prod_{k=1}^{n}\Big{(}\frac{\delta}{\text{i}\delta J_{a_{k}}}\Big{)}W[J]\Big{|}_{J=0}\;,$ (73)
where, to alleviate notations, we employ from now on de Witt convention, whereby the roman letter indices stand for the field components and their respective spacetime arguments. In particular, the two-point connected Green’s functions $G_{ab}^{(2)}$ are given by
$G^{(2)}_{ab}\equiv-W^{(2)}_{ab}[J=0]=\Big{(}\frac{\delta^{2}}{\text{i}\delta J_{a}\text{i}\delta J_{b}}\Big{)}W[J]\Big{|}_{J=0}\;.$ (74)
They can be gathered in a $4\times 4$ matrix as
$G^{(2)}=\begin{pmatrix}{\mathbb{G}^{K}}&&{\mathbb{G}^{R}}\\ {\mathbb{G}^{A}}&&{\mathbb{0}}_{2\times 2}\end{pmatrix}\;,$ (75)
where the $2\times 2$ matrices ${\mathbb{G}}$ are named Keldysh ${\mathbb{G}^{K}}$, advanced ${\mathbb{G}^{A}}$ and retarded ${\mathbb{G}^{R}}$ propagators, and are constrained by causality as ${\mathbb{G}^{A}}(\omega)=({\mathbb{G}^{R}}(\omega))^{*}$.
Finally, we define the effective action $\Gamma$ as the Legendre-Fenchel transform of $W$ w.r.t the one-point expectation value of the fields $\phi\equiv\langle\varphi\rangle_{J}$
$\Gamma[\phi]\equiv\inf_{J[\phi]}\big{(}W[J]-\sum_{a}J_{a}\phi_{a}\big{)}\;,$ (76)
where the sum stands for summation over field and source components, as well as summation/integration over their spacetime arguments. For a system at equilibrium, the effective action corresponds to the Gibbs free energy and the correlation functions obtained by taking functional derivatives of $\Gamma$ with respect to $\phi$
$\Gamma^{(n)}_{a_{1}\dots a_{n}}[\phi]\equiv\frac{\delta^{n}}{\delta\phi_{a_{1}}\dots\delta\phi_{a_{n}}}\Gamma[\phi]\;,$ (77)
correspond to the renormalised vertex functions of the theory. From the Legendre relation, one deduces that the matrix of second derivatives of the effective action is the inverse of the propagator matrix $G^{(2)}$
$G^{(2)}=\Big{[}\Gamma^{(2)}\big{|}_{\phi=0}\Big{]}^{-1}\;.$ (78)
In order to obtain the self-energy, following usual field-theoretical methods, we compute $\Gamma^{(2)}$ within a one-loop expansion at small interaction. At one-loop order, it is given by
$\Gamma_{\text{(1-loop)}}=S-\frac{\text{i}}{2}\operatorname{Tr}\ln S^{(2)}\;,$ (79)
where the trace means again summation over the fields and summation/integration over their spacetime arguments. Taking two functional derivatives of this expression with respect to the fields, one obtains
$\Big{[}\Gamma^{(2)}_{ab}\Big{]}_{\text{(1-loop)}}=G^{-1}_{0,ab}-\frac{\text{i}}{2}\frac{\delta^{2}}{\delta\phi_{a}\delta\phi_{b}}\operatorname{Tr}\ln S^{(2)}\;.$ (80)
Identifying with the definition (20) for the self-energy yields
$\Big{[}\Sigma_{ab}\Big{]}_{\text{(1-loop)}}=\frac{\text{i}}{2}\frac{\delta^{2}}{\delta\phi_{a}\delta\phi_{b}}\operatorname{Tr}\ln S^{(2)}\Big{|}_{\phi=0}\;.$ (81)
This can be represented in terms of Feynman diagrams. With a four-point vertex $(-U)$ and propagator lines corresponding to the Green’s function components, the only diagrams one can construct are tadpole diagrams. Because of the integral over the frequency in the loop, the only non-vanishing contributions stem from diagrams constructed with the Keldysh propagator. This is a consequence of causality: The advanced and retarded propagators only have poles in a half complex plane and vanish upon integration on frequency. One finds
$\Sigma_{\text{(1-loop)}}=\begin{pmatrix}0&&0&&0&&\Sigma_{1}\\ 0&&0&&\Sigma_{1}&&0\\ 0&&\Sigma_{1}&&0&&0\\ \Sigma_{1}&&0&&0&&0\end{pmatrix}\;,$ (82)
where $\Sigma_{1}$ is given by
$\Sigma_{1}=\frac{\text{i}}{2}\times 2\times(-U)\int_{-\pi/a}^{\pi/a}a\frac{\text{d}p}{2\pi}\int\frac{\text{d}\omega}{2\pi}\frac{-\text{i}\gamma}{(\omega+2J\cos(pa))^{2}+\kappa_{0}^{2}}=\dfrac{\gamma U}{2\kappa_{0}}\,,$ (83)
where the factor $2$ takes the multiplicity of the diagram into account, the factor $-U$ arises from the $4$-point vertex, and the Keldysh propagator forming the loop is integrated over.
E.5 Time-reversal symmetry
A (quantum) system at equilibrium is invariant under time reversal. This definition of equilibrium manifests in the Keldysh action as an invariance under a discrete symmetry transformation [70]. Considering the entire quantum evolution, it is easy to verify that the action (4.1) does not fulfill the time reversal symmetry transformation, as the quadratic part in the quantum fields does not contain the term $\coth\beta\omega$, relating it to the Bose-Einstein distribution $n(\omega)=1/(\text{e}^{\beta\omega}\,-1)$.
Also, in the semi-classical limit, i.e. in the limit of large temperature $T$, this discrete transformation simplifies to
$\left\{\begin{array}[]{l l}\mathcal{T}\Phi_{c,j}(t)&=\sigma_{x}\Phi_{c,j}(-t)\\ \mathcal{T}\Phi_{q,j}(t)&=\sigma_{x}\Big{(}\Phi_{q,j}(-t)+\frac{\text{i}}{2T}\partial_{t}\Phi_{c,j}(-t)\Big{)}\,,\end{array}\right.$ (84)
where $\Phi_{\nu}$ denotes the vectors $\Phi_{\nu,j}(t)\equiv(\varphi_{\nu,j}(t),\varphi^{*}_{\nu,j}(t))^{T}$ with $\nu=c,q$. This transformation is equivalent to the well-known relation for classical dynamical equilibrium systems in the Martin-Siggia-Rose formalism [80, 81]. On can check that the action (4.1) is not invariant under this transformation. Indeed, the terms in the first line of (4.1) can only be time-reversal symmetric in the presence of the additional coherent contribution
$\int\text{d}t\sum_{j}\kappa\,\Big{(}\varphi_{q,j}^{*}\,\partial_{t}\varphi_{c,j}-\varphi_{c,j}^{*}\,\partial_{t}\varphi_{q,j}\Big{)}\,,\qquad\hbox{with}\;\;\kappa=-\frac{\gamma}{4T}$
and with $\gamma_{l}=\gamma_{p}$, i.e. $\kappa_{0}=0$. Hence, for $\kappa=0$ and $\kappa_{0}\neq 0$, the system breaks time-reversal symmetry, it is genuinely out-of-equilibrium.
Appendix F Decay rate at large interactions
@@
For $J=0$, each site decouples. Hence, in the strong interaction limit $U\rightarrow\infty$, one is left with solving the problem of $L$ independent dissipative Kerr resonators. This is a well-known exactly solvable problem [69]. One can exploit a weak symmetry to block-diagonalise the Lindbladian and then introduce an operator which renders the Lindbladian quadratic in the creation and annihilation superoperators, i.e. the operators acting on density matrices. Equivalently, one can introduce a time-dependent gauge transformation to obtain a quadratic action as shown in [59]. The Kerr non-linearity then transforms into a fluctuating frequency depending on the number density which in turn leads to dephasing. The result for the retarded Green’s function is given by [69]
@@
The last term is suppressed in the large $t$ limit. Expanding the second term for $U\gg\gamma_{l}$, we find that
@@
such that the renormalised decay for $U\gg\gamma_{l}$ reads
@@
Appendix G Details on the numerical calculations
G.1 Exact diagonalisation
Numerical results were obtained by exact diagonalisation of the Hamiltonian or Lindbladian in the Fock basis of particle occupation numbers per site and the time evolution implemented in the qutip library [82, 83]. In the case of the closed-system evolution, we additionally use the conservation of particle number, as explained in section G.5.
G.2 Fit of the decay
@@
Restricting the dimension of the local Hilbert space, we obtain results for all interaction strength by exact diagonalisation. Using this method, we extract the slope of ${\cal G}(t)$ by a linear fit to the oscillation curve. An example can be found in Fig. 10. Due to the oscillations on different time scales – note that we vary both $U$ and $J$ over several orders of magnitude –, the result for $\kappa$ can depend on the time intervals chosen. In turn, fitting a family of 20 curves in different time intervals gives an estimate of the variance of the fitted decay, which is the dominant error source of the fit.
@@
Figure 10: Logarithm of the retarded Green’s function ${\cal G}(t)=-\Re\Big{[}\ln\text{i}G_{00}^{R}(t,0)\Big{]}$ (dimensionless) obtained from exact diagonalisation (red solid line), as a function of time (in units of $\tau_{l}$) for the site $j=0$. Within the single-pole approximation, the decay rate is estimated by fitting ${\cal G}(t)$ by a linear model. The result of the fit is shown as the black solid line. The parameters are $U=5.623\gamma_{l}$, $J=\gamma_{l}$, $z=0.1$, $L=4$, $N_{s}=4$.
G.3 Effect of the Hilbert space truncation
@@
In the main text we showed that in the case $J=0$ we recover the analytically predicted renormalisation of the decay $\kappa(U)$. In order to benchmark the numerical truncation over the local Hilbert space, we here investigate the dependence of $\kappa(U)$ for large hopping values, taking $J/\gamma_{l}=10$. We show in Fig. 11 the analytical (single-site) prediction, holding in the limit of vanishing $J$, and the numerical results for $N_{s}\in\{3,4,5\}$. For large $U$, we expect a deviation from the single-site result and we observe that the reached plateau is independent of the local Hilbert space truncation. In the weak interaction limit, we interpret the closeness of the results for $N_{s}=4$, $N_{s}=5$ with the exact prediction as an indication of numerical convergence.
@@
Figure 11: Decay $\kappa(U)$ for $J=10\gamma_{l}$, $z=0.1$, and $L=4$, and for different local truncation of the Hilbert space $N_{s}$ in the numerics. We expect to recover the non-interacting solution for small $U$ and a deviation from the single-site solution in the limit of large interaction.
G.4 System size dependence
@@
We further investigate the dependence of the reached plateau value of the decay for large $U$ and $J$ on the system size $L$. From Fig. 11, we conclude that a truncation of $N_{s}=3$ is sufficient at large $U$ to converge numerically at the correct plateau value. Hence, we fix $N_{s}=3$ and vary the system size $L$. The dependence of the renormalised decay $\kappa$ on $L$ is shown in Fig. 12, in a limit of large hopping and interaction (numerical values: $J/\gamma_{l}=10$ and $U/\gamma_{l}=100$). We observe that the effective decay increases monotonously with the system size for these small system sizes considered.
@@
Figure 12: Plateau value of the decay $\kappa_{\infty}$ for large interaction $U=100\gamma_{l}$ and hopping $J=10\gamma_{l}$ at different system sizes $L=\{1,2,4,6\}$. The parameters are $z=0.1$, and $N_{s}=3$.
G.5 Exact diagonalisation for the unitary system
@@
We calculate the Green’s function as the overlap of two time-evolved states. Since we are interested in the unitary time evolution, the dimensionality of the problem greatly reduces, and is further diminished by the symmetries present in the problem (particle number conservation, momentum conservation and spatial parity conservation). We calculate the Green’s function as
@@
G.6 Simulations with matrix product states
In order to produce the plots in Fig. 6 we employ a representation of the many-body state as a matrix-product state [84] and code our algorithm using the Itensors library [85, 86]. We prepare the ground state considering a maximum number of bosons per site equal to $2$ on a lattice with periodic boundary condition and at density $N/L=1/4$. We subsequently compute the lesser and greater Green’s functions $G^{<}_{0j}(t)=-i\langle b_{j}^{\dagger}(t)b_{0}\rangle$ and $G^{>}_{0j}(t)=-i\langle b_{0}b_{j}^{\dagger}(t)\rangle$ and using these data we reconstruct the spectral functions plotted in the figure. For the time evolution, we fix $J=1$ and use a time step of $\delta t=0.01$. The chemical potential is obtained from the ground state energy $E_{N}$ of a $N$-particle system by taking the discrete derivative $\mu=E_{N+1}-E_{N}$, where $E_{N}$ is obtained from the DMRG calculations.